# Towards Fast Coarse-graining and Equation Discovery with Foundation Inference Models

### v1

@@ 1 Introduction
We propose instead to decouple the two problems by relying on the recently introduced Foundation Inference Models (FIMs). FIMs are pretrained models that perform in-context (or zero-shot) estimation of the infinitesimal generators of dynamical systems, and that fall under the umbrella of simulation-based inference models [Cranmer et al., 2020]. Examples include FIMs for MJPs [Berghaus et al., 2024], ODEs [d’Ascoli et al., 2024, Seifner et al., 2024], SDEs [Seifner et al., 2025], and point processes [Berghaus et al., 2025].
@@ Acknowledgments
This research has been funded by the Federal Ministry of Education and Research of Germany and the state of North-Rhine Westphalia as part of the Lamarr-Institute for Machine Learning and Artificial Intelligence.

### v2

@@ 1 Introduction
We propose instead to decouple the two problems by relying on the recently introduced Foundation Inference Models (FIMs). FIMs are pretrained models that perform in-context (or zero-shot) estimation of the infinitesimal generators of dynamical systems, and that fall under the umbrella of simulation-based inference models [Cranmer et al., 2020]. Examples include FIMs for MJPs [Berghaus et al., 2024], ODEs [d’Ascoli et al., 2024, Seifner et al., 2024, Mauel et al., 2025], SDEs [Seifner et al., 2025], and point processes [Berghaus et al., 2025a, b].
@@ Acknowledgments
This research has been funded by the Federal Ministry of Education and Research of Germany and the state of North-Rhine Westphalia as part of the Lamarr Institute for Machine Learning and Artificial Intelligence.